A large pet-food manufacturer is considering buying a small boutique cat food business to add to their portfolio of pet foods. The head of the finance division has approached you to conduct a financial analysis to determine the most the firm should pay for the cat food business. The valuation of the cat-food business is based on cash flows of $180,500 per year over a five year period. The target business has the same risk as the firm’s overall operations. The cost of equity is 15 percent and the cost of debt is 3 percent on an after-tax basis. The firm’s capital structure consists of 10 million in equity and 8 million in debt. What is the most the firm can pay for the business per its required return (WACC) ?
Answer #1

Cost of equity = re = 15% and after tax cost of debt = rd = 3%

Total debt = D = $8 million, Total Equity = E = $10 million

Total capital = V = D + E = 8 + 10 = $18 million

Weight of debt in total capital = D/V = 8/18, Weight of Equity in total capital = E/V = 10/18

WACC = (D/V)(rd) + (E/V)(re) = (8/18)(3%) + (10/18)(15%) = 44.44% x 3% + 55.56% x 15% = 1.3332% + 8.3340% = 9.6672% = 9.67% (rounded to two decimal places)

Cash flow per year = $180500, No of years = 5

Value of Cat Food business = Sum of present values of cash flows discounted at WACC

= CF1 / (1 + WACC) + CF2 / (1 + WACC)2 + CF3 / (1 + WACC)3 + CF4 / (1 + WACC)4 + CF5 / (1 + WACC)5

= 180500 / (1 + 9.67%) + 180500 / (1 + 9.67%)2 + 180500 / (1 + 9.67%)3 + 180500 / (1 + 9.67%)4 + 180500 / (1 + 9.67%)5

= 164584.6630 + 150072.6389 + 136840.1923 + 124774.4983 + 113772.6801

= 690044.6726 = 690044.67

Most the firm can pay for its business = $690044.67